OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 2 to the fifth power, over 8 = 22 By finding the quotient of the bases to be,one fourthand cancelling common factors By finding the quotient of the bases to be, one fourthand simplifying the expression By simplifying 8 to 23 to make both powers base two, and subtracting the exponents By simplifying 8 to 23 to make both powers base two, and adding the exponents

OpenStudy (anonymous):

please help

OpenStudy (solomonzelman):

2^5 / 8 = 2^5 / 2^3 = 2^(5-3)=2^2

OpenStudy (anonymous):

3rd

OpenStudy (anonymous):

thanks guys i owe you